# High-low analysis and small cap decoupling over non-Archimedean local fields

### v1

@@ 2.3 Broad/narrow analysis
For each cube $B_{R}$ of sidelength $R$ and each $\alpha>0$,
$\alpha^{4}\mu\Big{(}\{x\in B_{R}:|f(x)|>\alpha\}\Big{)}\leq 12(1+N^{-5})^{10}\mathbf{p}^{12}N^{10}\sum_{0\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
(#$(\#\tau_{k}\prec\tau_{k-1})\max_{\tau_{k}\neq\tau_{k}^{\prime}}|f_{m,\tau_{k}}^{\mathcal{B}}(x)|>\left(1-\left(1+\frac{1}{N-1}\right)^{-1}\right)|f_{m,\tau_{k-1}}^{\mathcal{B}}(x)|.$
@@
By the pruning inequalities, for each $\theta$ we may take $|f_{m,\theta}^{\mathcal{B}}|\leq|f_{N,\theta}|$. By the definition of the pruning, for each $\theta$,
@@
$\sum_{\tau_{k-1}}\int_{\mathrm{Broad}_{k,m}(\tau_{k-1})}|f_{m,\tau_{k-1}}^{\mathcal{B}}|^{4}\leq 12\mathbf{p}^{12}N^{5}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}$
@@
$\int_{\mathrm{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})}|f_{m,\tau_{k}}^{\mathcal{B}}f_{m,\tau_{k}^{\prime}}^{\mathcal{B}}|^{2}\leq 4\mathbf{p}^{2}\int_{\mathbb{Q}_{p}^{2}}\left|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2},$
where $\ell=\max(m-1,k)$. By Prop. (2.3),
$\int_{\mathbb{Q}_{p}^{2}}\left|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2}\leq 3\mathbf{p}^{4}N\sum_{m\leq s\leq N}\sum_{\tau_{s}\prec\tau_{k-1}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$\sum_{\tau_{k-1}}\int_{\mathrm{Broad}_{k,m}(\tau_{k-1})}|f_{m,\tau_{k-1}}^{\mathcal{B}}|^{4}\leq 12\mathbf{p}^{12}N^{5}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$\alpha^{4}\mu(U_{\alpha}^{m})\leq N^{4}(1+N^{-5})^{10}(1-(2^{1/2}eN)^{-1})^{-1}\int_{U_{\alpha}^{m}}|f_{m}^{\mathcal{B}}|^{4}.$
@@
$\sum_{\tau_{k-1}}\int_{\mathrm{Broad}_{k,m}(\tau_{k-1})}|f_{m,\tau_{k-1}}^{\mathcal{B}}|^{4}\leq 12\mathbf{p}^{12}N^{5}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$\alpha^{4}\mu(U_{\alpha}^{m})\leq 12(1+N^{-5})^{10}(2^{1/2}eN)^{-1})^{-1}\mathbf{p}^{12}N^{9}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2},$
and hence
$\alpha^{4}\mu(U_{\alpha})\leq 13(1+N^{-5})^{10}(1-(2^{1/2}eN)^{-1})^{-1}\mathbf{p}^{12}N^{10}\sum_{0\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2},$ (2.6)
@@ 2.4 Reduction to local estimates
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|1_{Q_{R}}f|>\alpha\})\leq 13\frac{(1+N^{-5/2})^{10}}{1-(2^{1/2}eN)^{-1}}N^{10}\mathbf{p}^{12}\sum_{R^{-1/2}\leq s\leq 1}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|(1_{Q_{R}}f)_{\theta}|^{2}\right)^{2}.$
@@
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|f(x)|>\alpha\})\leq 13\frac{(1+N^{-5/2})^{10}}{1-(2^{1/2}eN)^{-1}}N^{10}\mathbf{p}^{12}\sum_{R^{-1/2}\leq s\leq 1}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\sum_{Q_{R}}\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|(1_{Q_{R}}f)_{\theta}|^{2}\right)^{2}.$
@@
Theorem (1.1) will be proved via first establishing the critical case $(p,q)=(2+\frac{2}{\beta},\frac{2+2\beta^{-1}}{2\beta^{-1}-1})$, and then interpolating with the easier endpoints $(\infty,1)$ and $(3,\infty)$, together with Hölder inequalities. We will repeatedly cite the result of Theorem (1.2), and we will abbreviate the constants as
@@
Suppose $p\geq 4$, $\lambda>0$, and $\mathfrak{C}_{p}>1$. Let $0\leq k\leq N$ be arbitrary, and fix a canonical scale cap $\tau_{k}$. Suppose as before that $\Gamma_{\beta}(R^{-1})$ is a partition of $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$ into approximate $R^{-\beta}\times R^{-1}$ boxes $\gamma$. Assume $f=\sum_{\gamma}f_{\gamma}$ satisfies the following regularity properties.
@@
    $\|f_{\gamma}\|_{2}^{2}\leq\lambda^{2-p}\mathfrak{C}_{p}\|f_{\gamma}\|_{p}^{p}$ for each $\gamma$ and each $p\geq 1$.
@@
$\begin{split}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}&\leq\mathfrak{C}_{p}\big{[}2^{\frac{1}{2}}eN(\#\tau_{k})\alpha^{-1}\big{]}^{p-4}\\ &\times\left(\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\times\#(\gamma\prec\tau_{k})\right)^{\frac{p}{2}-1}\sum_{\gamma\prec\tau_{k}}\|f_{\gamma}\|_{p}^{p}.\end{split}$ (3.1)
@@
Multiplying the left-hand side of (3.1) by the $(\frac{p}{2}-2)$-power of the latter display, we obtain the estimate
$\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}\leq\big{[}2^{1/2}eN(\#\tau_{k})\alpha^{-1}\big{]}^{p-4}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\gamma_{k}\prec\tau_{k}}|f_{\gamma_{k}}|^{2}\right)^{\frac{p}{2}}.$ (3.2)
@@
$\begin{split}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}\leq&\,(\#\tau_{k})^{p-4}\lambda^{p-2}\alpha^{4-p}\left(\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\times\#(\gamma\prec\tau_{k})\right)^{\frac{p}{2}-1}\sum_{U\in\mathcal{G}_{\tau_{k}}}\int_{U}\sum_{\gamma_{k}\prec\tau_{k}}|f_{\gamma_{k}}|^{2},\end{split}$
@@
$\sum_{U\in\mathcal{G}_{\tau_{k}}}\int_{U}\sum_{\gamma_{k}\prec\tau_{k}}|f_{\gamma_{k}}|^{2}=\int\sum_{\gamma\prec\tau_{k}}|f_{\gamma}|^{2}\leq\lambda^{2-p}\mathfrak{C}_{p}\sum_{\gamma\prec\tau_{k}}\|f_{\gamma}\|_{p}^{p}.$
@@
$\begin{split}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}&\leq\mathfrak{C}_{p}\big{[}2^{1/2}eN(\#\tau_{k})\alpha^{-1}\big{]}^{p-4}\\ &\times\left(\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\times\#(\gamma\prec\tau_{k})\right)^{\frac{p}{2}-1}\sum_{\gamma\prec\tau_{k}}\|f_{\gamma}\|_{p}^{p},\end{split}$
@@
Let $f$ be Schwartz-Bruhat with Fourier support in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$, such that $\max_{\theta}\|f_{\theta}\|_{\infty}=1$. Let $(p,q)=(2+\frac{2}{\beta},\frac{2+2\beta^{-1}}{2\beta^{-1}-1})$. Then, for each $R^{-1/2}\leq\alpha\leq R^{1/2}$, we have the estimate
$\alpha^{p}\mu\big{(}\{x:|f(x)|>\alpha\}\big{)}|\lesssim\mathbf{p}^{12}N^{2p-2}(\log R)^{p+2}\max\big{(}R^{\beta(p-\frac{p}{q}-1)-1}R^{\beta(\frac{p}{2}-\frac{p}{q})}\big{)}\Big{(}\sum_{\gamma}\|f_{\gamma}\|_{p}^{q}\Big{)}^{\frac{p}{q}}.$ (3.3)
@@
where the $\lambda$ range over values of the form $\lambda_{k+1}=\lfloor e^{-1}\lambda_{k}\rfloor$, $\lambda_{0}=1$, and $\eta$ is Schwartz-Bruhat, supported on $B_{R}$, and uniformly bounded by $1$. We abbreviate
@@
$e^{-p}(\#\mathcal{U}_{\lambda})R^{1+\beta}\lambda^{p}\leq\Big{\|}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big{\|}_{p}^{p}\leq(\#\mathcal{U}_{\lambda})R^{1+\beta}\lambda^{p},$
@@
$\Big{\|}\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big{\|}_{2}^{2}\leq(\#\mathcal{U}_{\lambda})R^{1+\beta}\leq e^{p}\Big{\|}\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big{\|}_{p}^{p}$ (3.4)
@@
$\Big{\|}\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big{\|}_{p}^{p}\leq(\#\mathcal{U}_{\lambda}^{\gamma})R^{1+\beta}\leq e^{2}\Big{\|}\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big{\|}_{2}^{2}.$ (3.5)
For each $1\leq\mathfrak{r}\leq R$ and each $\lambda$, write $\Gamma_{\beta}^{\lambda;\mathfrak{r}}(R^{-1})$ to be the collection of $\gamma\in\Gamma_{\beta}^{\lambda}(R^{-1})$ such that #$\#\mathcal{U}_{\lambda}^{\gamma}\in(e^{-1}\mathfrak{r},\mathfrak{r}]$. Define for $\gamma\in\Gamma_{\beta}^{\lambda}(R^{-1})$
@@
$g^{(\lambda,\mathfrak{r})}=\sum_{\gamma\in\Gamma_{\beta}^{\lambda;\mathfrak{r}}(R^{-1})}g_{\gamma}^{(\lambda)}.$
We may observe that $g^{(\lambda,\mathfrak{r})}$ is a Schwartz-Bruhat function with Fourier support in the $R^{-1}$-neighborhood of the truncated parabola. Thus, for each $\lambda,\mathfrak{r}$, and $\mathfrak{a}>0$ we have
$\mathfrak{a}^{4}\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}\leq CN^{E_{2}}\sum_{0\leq k\leq N}\sum_{\tau_{k}}\sum_{U\in\mathcal{G}_{\tau_{k}}[\lambda g^{(\lambda,\mathfrak{r})};\mathfrak{a}]}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|\lambda g_{\theta}^{(\lambda,\mathfrak{r})}|^{2}\right)^{2},$
where we have written $\mathcal{G}_{\tau_{k}}[\lambda g^{(\lambda,\mathfrak{r})};\mathfrak{a}]$ to record that the pruning is with respect to the function $\lambda g^{(\lambda,\mathfrak{r})}$ with the amplitude parameter $\mathfrak{a}$. For each $0\leq k\leq N$ and each $1\leq\mathfrak{s}\leq R^{1/2}$, let $\mathcal{T}_{k}(\mathfrak{s})$ denote the collection of $\tau_{k}$ such that #$\#\{\gamma\prec\tau_{k}:g_{\gamma}^{(\lambda,\mathfrak{r})}\neq 0\}\in(e^{-1}\mathfrak{s},\mathfrak{s}]$. By pigeonholing, for each $k$ we may find $\mathfrak{s}^{k}$ such that
$\mathfrak{a}^{4}\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}\leq 2^{-1}CN^{E_{2}}(\log R)\sum_{0\leq k\leq N}\sum_{\tau_{k}\in\mathcal{T}_{k}(\mathfrak{s}^{k})}\sum_{U\in\mathcal{G}_{\tau_{k}}[\lambda g^{(\lambda,\mathfrak{r})};\mathfrak{a}]}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|\lambda g_{\theta}^{(\lambda,\mathfrak{r})}|^{2}\right)^{2}.$

### v2

@@ 2.3 Broad/narrow analysis
Suppose $N>6$. For each cube $B_{R}$ of sidelength $R$ and each $\alpha>0$,
$\alpha^{4}\mu\Big(\{x\in B_{R}:|f(x)|>\alpha\}\Big)\leq 20\mathbf{q}^{-3}\mathbf{p}^{10}N^{10}\sum_{0\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$(\#\tau_{k}\prec\tau_{k-1})\max_{\tau_{k}\neq\tau_{k}^{\prime}}|f_{m,\tau_{k}}^{\mathcal{B}}(x)|>\left(1-\left(1+\frac{1}{N-1}\right)^{-1}\right)|f_{m,\tau_{k-1}}^{\mathcal{B}}(x)|.$
@@
By the pruning inequalities (2.1), for each $\theta$ we may take $|f_{m,\theta}^{\mathcal{B}}|\leq|f_{N,\theta}|$. By the definition of the pruning, for each $\theta$,
@@
$\sum_{\tau_{k-1}}\int_{\mathrm{Broad}_{k,m}(\tau_{k-1})}|f_{m,\tau_{k-1}}^{\mathcal{B}}|^{4}\leq 18\mathbf{q}^{-3}\mathbf{p}^{10}N^{5}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}$
@@
$\int_{\mathrm{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})}|f_{m,\tau_{k}}^{\mathcal{B}}f_{m,\tau_{k}^{\prime}}^{\mathcal{B}}|^{2}\leq 6\mathbf{q}^{-2}\mathbf{p}^{2}\int_{\mathbb{Q}_{p}^{2}}\left|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2},$
where $\ell=\max(m-1,k)$. By Prop. (2.3) and the wave envelope expansion lemma,
$\int_{\mathbb{Q}_{p}^{2}}\left|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2}\leq 3\mathbf{q}^{-1}\mathbf{p}^{4}N\sum_{s=\ell+1}^{N}\sum_{\tau_{s}\prec\tau_{k-1}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$\sum_{\tau_{k-1}}\int_{\mathrm{Broad}_{k,m}(\tau_{k-1})}|f_{m,\tau_{k-1}}^{\mathcal{B}}|^{4}\leq 18\mathbf{q}^{-3}\mathbf{p}^{10}N^{5}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$\alpha^{4}\mu(U_{\alpha}^{m})\leq N^{4}(1+N^{-4})^{10}(1-(2^{1/2}eN)^{-1})^{-1}\int_{U_{\alpha}^{m}}|f_{m}^{\mathcal{B}}|^{4}.$
@@
$\sum_{\tau_{k-1}}\int_{\mathrm{Broad}_{k,m}(\tau_{k-1})}|f_{m,\tau_{k-1}}^{\mathcal{B}}|^{4}\leq 18\mathbf{q}^{-3}\mathbf{p}^{10}N^{5}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2}.$
@@
$\alpha^{4}\mu(U_{\alpha}^{m})\leq 18(1+N^{-4})^{10}(2^{1/2}eN)^{-1})^{-1}\mathbf{p}^{10}N^{9}\sum_{m\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2},$
and hence, using the $m=0$ estimate and trivial estimates using $N>6$, we conclude that
$\alpha^{4}\mu(U_{\alpha})\leq 20\mathbf{q}^{-3}\mathbf{p}^{10}N^{10}\sum_{0\leq s\leq N}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|f_{\theta}|^{2}\right)^{2},$ (2.7)
@@ 2.4 Reduction to local estimates
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|1_{Q_{R}}f|>\alpha\})\leq 20\mathbf{q}^{-3}\mathbf{p}^{10}N^{10}\sum_{R^{-1/2}\leq s\leq 1}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|(1_{Q_{R}}f)_{\theta}|^{2}\right)^{2}.$
@@
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|f(x)|>\alpha\})\leq 20\mathbf{q}^{-3}\mathbf{p}^{10}N^{10}\sum_{R^{-1/2}\leq s\leq 1}\sum_{\tau_{s}}\sum_{U\in\mathcal{G}_{\tau_{s}}}\mu(U)\sum_{Q_{R}}\left(\fint_{U}\sum_{\theta\prec\tau_{s}}|(1_{Q_{R}}f)_{\theta}|^{2}\right)^{2}.$
@@
Theorem (1.1) will be proved via first establishing the critical case $(r,q)=(2+\frac{2}{\beta},\frac{2+2\beta^{-1}}{2\beta^{-1}-1})$, and then interpolating with the easier endpoints $(\infty,1)$ and $(3,\infty)$, together with Hölder inequalities. We will repeatedly cite the result of Theorem (1.2), and we will abbreviate the constants as
@@
Suppose $r\geq 4$, $\lambda>0$, and $\mathfrak{C}_{r}>1$. Let $0\leq k\leq N$ be arbitrary, and fix a canonical scale cap $\tau_{k}$. Suppose as before that $\Gamma_{\beta}(R^{-1})$ is a partition of $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$ into approximate $R^{-\beta}\times R^{-1}$ boxes $\gamma$. Assume $f=\sum_{\gamma}f_{\gamma}$ satisfies the following regularity properties.
@@
    $\|f_{\gamma}\|_{2}^{2}\leq\lambda^{2-r}\mathfrak{C}_{r}\|f_{\gamma}\|_{r}^{r}$ for each $\gamma$ and each $r\geq 1$.
@@
$\begin{split}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}&\leq\mathfrak{C}_{r}\big[2^{\frac{1}{2}}eN(\#\tau_{k})\alpha^{-1}\big]^{r-4}\\ &\times\left(\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\times\#(\gamma\prec\tau_{k})\right)^{\frac{r}{2}-1}\sum_{\gamma\prec\tau_{k}}\|f_{\gamma}\|_{r}^{r}.\end{split}$ (3.1)
@@
Multiplying the left-hand side of (3.1) by the $(\frac{r}{2}-2)$-power of the latter display, we obtain the estimate
$\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}\leq\big[2^{1/2}eN(\#\tau_{k})\alpha^{-1}\big]^{r-4}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\gamma_{k}\prec\tau_{k}}|f_{\gamma_{k}}|^{2}\right)^{\frac{r}{2}}.$ (3.2)
@@
$\begin{split}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}&\leq 2^{\frac{r}{2}-2}e^{r-4}N^{r-4}(\#\tau_{k})^{r-4}\lambda^{r-2}\alpha^{4-r}\\ &\times\left(\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\times\#(\gamma\prec\tau_{k})\right)^{\frac{r}{2}-1}\sum_{U\in\mathcal{G}_{\tau_{k}}}\int_{U}\sum_{\gamma_{k}\prec\tau_{k}}|f_{\gamma_{k}}|^{2},\end{split}$
@@
$\sum_{U\in\mathcal{G}_{\tau_{k}}}\int_{U}\sum_{\gamma_{k}\prec\tau_{k}}|f_{\gamma_{k}}|^{2}=\int\sum_{\gamma\prec\tau_{k}}|f_{\gamma}|^{2}\leq\lambda^{2-r}\mathfrak{C}_{r}\sum_{\gamma\prec\tau_{k}}\|f_{\gamma}\|_{r}^{r}.$
@@
$\begin{split}\sum_{U\in\mathcal{G}_{\tau_{k}}}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}\right)^{2}&\leq\mathfrak{C}_{r}\big[2^{1/2}eN(\#\tau_{k})\alpha^{-1}\big]^{r-4}\\ &\times\left(\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\times\#(\gamma\prec\tau_{k})\right)^{\frac{r}{2}-1}\sum_{\gamma\prec\tau_{k}}\|f_{\gamma}\|_{r}^{r},\end{split}$
@@
Suppose $R\geq 16$ and $N\geq 10$. Let $f$ be Schwartz–Bruhat supported in $B_{R}$ with Fourier support in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$, such that $\max_{\theta}\|f_{\theta}\|_{\infty}=1$. Let $(r,q)=(2+\frac{2}{\beta},\frac{2+2\beta^{-1}}{2\beta^{-1}-1})$. Then, for each $R^{-1/2}\leq\alpha\leq R^{1/2}$, we have the estimate
$\begin{split}\alpha^{r}\mu\big(\{x:|f(x)|>\alpha\}\big)|&\leq\frac{e^{4r-3}}{50\cdot 2^{3r/2}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{r+7}}(\log R)^{3r+10}\\ &\times\max\big(R^{\beta(r-\frac{r}{q}-1)-1},R^{\beta(\frac{r}{2}-\frac{r}{q})}\big)\Big(\sum_{\gamma}\|f_{\gamma}\|_{r}^{q}\Big)^{\frac{r}{q}}.\end{split}$ (3.3)
@@
where the $\lambda$ range over values of the form $\lambda_{k+1}=\lfloor e^{-1}\lambda_{k}\rfloor$, $\lambda_{0}=1$, and $\eta$ is Schwartz–Bruhat, supported on $B_{R}$, and uniformly bounded by $1$. We abbreviate
@@
$e^{-r}(\#\mathcal{U}_{\lambda})R^{1+\beta}\lambda^{r}\leq\Big\|\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big\|_{r}^{r}\leq(\#\mathcal{U}_{\lambda})R^{1+\beta}\lambda^{r},$
@@
$\Big\|\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big\|_{2}^{2}\leq(\#\mathcal{U}_{\lambda})R^{1+\beta}\leq e^{r}\Big\|\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big\|_{r}^{r}$ (3.4)
@@
$\Big\|\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big\|_{r}^{r}\leq(\#\mathcal{U}_{\lambda}^{\gamma})R^{1+\beta}\leq e^{2}\Big\|\frac{1}{\lambda}\sum_{U\in\mathcal{U}_{\lambda}}1_{U}f_{\gamma}\Big\|_{2}^{2}.$ (3.5)
For each $1\leq\mathfrak{t}\leq R$ and each $\lambda$, write $\Gamma_{\beta}^{\lambda;\mathfrak{t}}(R^{-1})$ to be the collection of $\gamma\in\Gamma_{\beta}^{\lambda}(R^{-1})$ such that $\mathcal{U}_{\lambda}^{\gamma}\in(e^{-1}\mathfrak{t},\mathfrak{t}]$. Define for $\gamma\in\Gamma_{\beta}^{\lambda}(R^{-1})$
@@
$g^{(\lambda,\mathfrak{t})}=\sum_{\gamma\in\Gamma_{\beta}^{\lambda;\mathfrak{t}}(R^{-1})}g_{\gamma}^{(\lambda)}.$
We may observe that $g^{(\lambda,\mathfrak{t})}$ is a Schwartz–Bruhat function with Fourier support in the $R^{-1}$-neighborhood of the truncated parabola. Thus, for each $\lambda,\mathfrak{t}$, and $\mathfrak{a}>0$ we have
$\mathfrak{a}^{4}\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)\leq CN^{E_{2}}\sum_{0\leq k\leq N}\sum_{\tau_{k}}\sum_{U\in\mathcal{G}_{\tau_{k}}[\lambda g^{(\lambda,\mathfrak{t})};\mathfrak{a}]}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|\lambda g_{\theta}^{(\lambda,\mathfrak{t})}|^{2}\right)^{2},$
where we have written $\mathcal{G}_{\tau_{k}}[\lambda g^{(\lambda,\mathfrak{t})};\mathfrak{a}]$ to record that the pruning is with respect to the function $\lambda g^{(\lambda,\mathfrak{t})}$ with the amplitude parameter $\mathfrak{a}$. For each $0\leq k\leq N$ and each $1\leq\mathfrak{s}\leq R^{1/2}$, let $\mathcal{T}_{k}(\mathfrak{s})$ denote the collection of $\tau_{k}$ such that $\{\gamma\prec\tau_{k}:g_{\gamma}^{(\lambda,\mathfrak{t})}\neq 0\}\in(e^{-1}\mathfrak{s},\mathfrak{s}]$. By pigeonholing, for each $k$ we may find $\mathfrak{s}^{k}$ such that
$\mathfrak{a}^{4}\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)\leq 2^{-1}CN^{E_{2}}(\log R)\sum_{0\leq k\leq N}\sum_{\tau_{k}\in\mathcal{T}_{k}(\mathfrak{s}^{k})}\sum_{U\in\mathcal{G}_{\tau_{k}}[\lambda g^{(\lambda,\mathfrak{t})};\mathfrak{a}]}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau_{k}}|\lambda g_{\theta}^{(\lambda,\mathfrak{t})}|^{2}\right)^{2}.$